OpenStudy (anonymous):

If g(x) = .5x + 3, find the value of g(0) + 2. A. 5 B. 4 C. 5.5 D. 7

jimthompson5910 (jim_thompson5910):

Replace x with 0 and evaluate to find g(0) Then add 2 to the result

OpenStudy (anonymous):

I dont get it :/

jimthompson5910 (jim_thompson5910):

g(x) = .5x + 3 g(0) = .5*0 + 3 <<--- Evaluate this to find g(0)

jimthompson5910 (jim_thompson5910):

What do you get?

OpenStudy (anonymous):

g(0)=3

jimthompson5910 (jim_thompson5910):

now add 2 to that to get your final answer

jimthompson5910 (jim_thompson5910):

g(0) = 3 g(0) + 2 = 3 + 2

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw